Odds converter

Convert an equivalent price between formats. The implied probability contains the bookmaker’s pricing assumptions and is not an independent forecast.

Worked example2.50 = +150 = 3/2; 1 / 2.50 = 40%.

Calculation

D = 1 + A/100 (A ≥ 100); D = 1 + 100/|A| (A ≤ −100); D = 1 + a/b.

Worked example

2.50 = +150 = 3/2; 1 / 2.50 = 40%.

American and decimal odds ↗

Useful next steps